zarkam21 (zarkam21):

Helpp please with: find the value of c that makes the polynomial a perfect square trinomial. x^2+8x+c

OpenStudy (hafeda):

@zarkam21 it is the same process u did just right now

zarkam21 (zarkam21):

I understand but I usually get assistance =)

zarkam21 (zarkam21):

Because I am new to the process

OpenStudy (hafeda):

DO the second term as follow : \[ \pm (\frac{ b }{ 2 } ) ^{2}\] it will be much better to do it first, u will learn more that way

zarkam21 (zarkam21):

so 8/2

zarkam21 (zarkam21):

(8/2)^2

OpenStudy (hafeda):

yes

zarkam21 (zarkam21):

16?

OpenStudy (hafeda):

yes now add 16 to both side of the equation , or add 16 and subtract 16

zarkam21 (zarkam21):

16x^2 +16+16?

OpenStudy (hafeda):

\[x ^{2} + 8x +16 - 16\]

OpenStudy (chiunlee12):

no, substitute 16 for 'c'

zarkam21 (zarkam21):

Oh okay I see what you did

OpenStudy (hafeda):

now u have a complete square and of \[(x+4) ^{2} - 16 \]

zarkam21 (zarkam21):

Oh okay and then c would be 16?

OpenStudy (chiunlee12):

hafeda, i think you are over solving the problem and yes, zarkam21, c would be 16 :D

OpenStudy (hafeda):

and that is constant out the square is the "C" @chiunlee12 i just want her to understand the process, and same time that ask it differently

OpenStudy (hafeda):

@zarkam21 yes c = 16

zarkam21 (zarkam21):

Would it be the same process for this

OpenStudy (chiunlee12):

@hafeda Yeah, I understand what you tried to do there, and your explanation was great

OpenStudy (chiunlee12):

Okay now that is almost the same as the first question

OpenStudy (danjs):

Wants a perfect square trinomial, a polynomial with 3 terms and factors like (a - b)^2 = (a - b)(a - b) = a^2 - 2ab + b^2 or (a + b)^2 = (a +b)(a + b) = a^2 + 2ab + b^2 notice the last term in these are squares, 2^2 ,3^2 , 4^2 , 6^2,.... x^2+8x+c to get an 8 in the middle from the sum of 2 *perfect square, you need to use 4 (x + 4)^2 = x^2 + 8x + 16

zarkam21 (zarkam21):

Okay and then we would continue to solve for b am I correct?

zarkam21 (zarkam21):

@3mar

OpenStudy (3mar):

3mar is here!

zarkam21 (zarkam21):

Yay =)

OpenStudy (3mar):

so? where did you reach?

OpenStudy (3mar):